# A Systems Thinking Approach to Algorithmic Fairness

### v4

@@ Abstract.
Systems thinking provides us with a way to model the algorithmic fairness problem by allowing us to encode prior knowledge and assumptions about where we believe bias might exist in the data generating process. We can then encode these beliefs as a series of causal graphs, enabling us to link AI/ML systems to politics and the law. This allows us to combine techniques from machine learning, causal inference, and system dynamics in order to capture different emergent aspects of the fairness problem. We can use systems thinking to help policymakers on both sides of the political aisle to understand the complex trade-offs that exist from different types of fairness policies, providing a sociotechnical foundation for designing AI policy that is aligned to their political agendas and with society’s values.
@@ 1. Introduction
For several years, AI researchers have struggled with how to build AI/ML systems that are fair and aligned to society’s values. Purely technical approaches that rely on correlation alone have fallen far short of what is needed to solve this difficult problem. Systems thinking offers a novel approach to the algorithmic fairness problem by allowing us to use causality to unify complex ideas from the social sciences and computer science into a comprehensive and integrated theory. By leveraging prior knowledge and assumptions about where we believe bias might exist in the data generating process, we can help policymakers on both sides of the aisle to design AI policy that is aligned to their political agendas.
@@
We will then turn our attention to the technical aspects of the fairness problem, which forms the bulk of this paper. We will use systems thinking to show how to combine analytical techniques from machine learning, causal inference, and system dynamics to model different facets of the fairness problem. Machine learning is necessary for building AI systems that can make fair predictions. Causal inference allows us to visualize different forms of bias, showing us what interventions are needed to make a machine learning model fair. System dynamics allows us to model the bias in the data generating process, providing counterintuitive explanations as to why disparities continue to persist across protected groups. Finally, we will demonstrate how these three techniques can form the basis of a new causal hierarchy for modeling algorithmic fairness.
@@ 2. Related work
Pearl has argued for a few decades that the discrimination problem is fundamentally a causality problem(Pearl, 2001). He has argued that fairness can be modeled using causal Bayesian networks (CBNs) as a series of direct, indirect, and spurious effects. This work was later expanded upon by Plecko and Bareinboim, who proposed the idea of a standard fairness model(Plecko and Bareinboim, 2022). They argued that there exists a causal template through which entire classes of fairness models can be derived using a small number of modeling assumptions. Kusner et al. argued for the use of counterfactuals in modeling fairness(Kusner et al., 2017). Chiappa and Isaac showed that certain paths within a CBN could be labeled as fair or unfair(Chiappa and Isaac, 2019). Another attempt of building causal fairness models was made by Nilforoshan et al., but their approach had encountered limitations(Nilforoshan et al., 2022). Barocas et al. devoted a chapter in their textbook on fairness and machine learning to discuss how causality might be used to model fairness(Barocas et al., 2023).
However, their causal approaches were not really sociotechnical and didn’t take full account of the social and political aspects of the fairness problem. Society has different ideological beliefs about what is fair and unfair based on different sets of assumptions about where bias is in the data generating process. Our paper addresses this issue by encoding society’s different mental models of bias into causal models using first principles. By taking into account a diversity of perspectives, we can understand the normative aspects of the fairness problem. This is necessary for linking machine learning to politics and ultimately the law.
Another key difference is that their causal approaches relied solely on causal inference, which has limitations in its ability to model the data generating process. Our approach leverages systems thinking and specifically system dynamics to gain a fuller understanding of the data generating process through the modeling of feedback loops. Martin has also explored modeling AI bias using community-based system dynamics(Jr, 2023, 2021). Furthermore, we link supervised machine learning to causal inference by recognizing that a CBN can be used as a higher level abstraction of a black box ML model. By bringing machine learning, causal inference, and systems thinking together, we can model the different aspects of the algorithmic fairness problem.
@@ 3. Towards a unified theory of fairness
To determine whether an outcome is biased and unfair, all we need to know is whether we would attribute a bad outcome as being internal or external. This makes algorithmic fairness a type of credit assignment problem. But since we usually focus on bad outcomes, this oftentimes becomes a blame assignment problem. That being said, identifying attribution is much harder in practice, as we will show later when we model fairness using system dynamics. Both internal and external factors may influence and reinforce each other, leading to vicious feedback loops. This becomes a particularly difficult problem when trying to explain disparities across protected groups like race, which we describe in greater detail in Appendix A.
This causality dilemma between attributing disparities to internal or external factors ultimately leads to the controversial and heated political debate over how best to reduce disparities across protected groups(Sowell, 2019; Donald, 2023; Mills, 1997; Kendi, 2017; Loury, 2002). Those on the political left are more likely to attribute disparities as being due to external factors, which provides the moral justification for strong government interventions like affirmative action or DEI. The left also tends to focus more on legal forms of discrimination like disparate impact or indirect discrimination. Those on the political right are more likely to attribute disparities as being due to internal factors, which provides the moral justification for limited government intervention like fairness through unawareness. The right also tends to focus more on legal forms of discrimination like disparate treatment or direct discrimination.
@@ 5.1. Mapping fairness to politics
Left-wing notions of fairness rely on affirmative action, which is associated with the criterion $D\perp\!\!\!\perp A$. That is, the decision $D$ is independent of the protected attribute $A$, which is achieved by using the protected attribute $A$ to correct for biases in the data $X$ that is used in the decision $D$. We imagine what the applicant’s decision would have been had they been a member of a counterfactual class. We then perform an intervention to obtain the same decision. For example, a college admissions decision would affirmatively use race to correct for deficiencies in the applicant’s merit due to structural discrimination. This leads to group fairness, where different groups receive more similar aggregate decision rates(Castelnovo et al., 2022; Verma and Rubin, 2018; Mehrabi et al., 2021; Binns, 2019). This form of fairness relies on Pearl’s first fundamental law of causal inference: counterfactuals and interventions.
@@ 5.2. Mapping discrimination to antidiscrimination law
When a protected attribute $A$ has a direct effect on a decision $D$ ($A\rightarrow D$), we might consider this as a form of disparate treatment or direct discrimination. This would map to overt discrimination (Figure 2b) and reverse discrimination (Figure 2f). To avoid disparate treatment or direct discrimination, the protected attribute $A$ should have no direct effect on the decision $D$.
@@
When a protected attribute $A$ has an indirect effect on a decision $D$, we might consider this as a form of disparate impact or indirect discrimination. This would map to covert discrimination (Figure 2d) and “no” discrimination (Figure 2h). Even though the protected attribute $A$ has no direct effect on a decision $D$ (making the policy neutral), the inclusion of biased data or exclusion of unbiased data could cause the protected attribute $A$ to have an indirect effect on the decision $D$ that adversely affects a protected group.
For disparate impact in the US, there is a three-step test(of Justice Civil Rights Division, 2024). The first step is to establish that a neutral policy has an adverse disparate impact on a protected group. This would involve testing whether there is a disparity in decisions $D$ across the protected class $A$. The second step is to establish the business necessity of that policy. This would involve arguing that the data $X$ being used has some causal relationship to the mediator $W$. The final step is to identify a less discriminatory alternative model, which would involve performing a model search.
CBNs provide an analytical framework to reason about whether a model is causing illegal discrimination. By being able to explicitly model the different forms of discrimination, CBNs provide a common language that businesses and regulators can use to ensure that machine learning models are compliant with laws and regulations. Appendix B describes the normative effects of the CBN from the different forms of fairness and discrimination.
@@ 6. Modeling fairness as a system dynamics problem
Finally, we add in the “shifting the burden / addiction” archetype. If one group is allocated less resources than another group, then it could breed resentment by the disadvantaged group. That group could then demand that the protected attribute $A$ be used to externally intervene on the allocation decision $D$. This external intervention is equivalent to fairness through affirmative action (Figure 2e). While it might reduce allocation disparities in the short term, it may actually create a dependence on the external intervention over the long term. This is because if allocation decisions $D$ are not based purely on data $X$, which is a measure of mediator $W$ (e.g. merit, qualifications), then the assignment of resources to the disadvantaged group may not lead to their success $Y$. For example, approving a loan to a minority applicant even though the applicant is not creditworthy could lead to a worse outcome (default) than if the loan was denied in the first place. That decreased success in outcome $Y$ for the less successful group could trigger a demand for an even stronger external intervention $A$ on the decision $D$, reinforcing this problem. In the meanwhile, attention may be taken away from the internal solution of optimizing for the mediator $W$ and the data $X$ so that the disadvantaged group would perform better under fairness through unawareness. This may especially be the case if the disadvantaged group believes that the mediator $W$ is biased and therefore does not focus on optimizing that metric. Appendix C describes the positive effects of the CLD from the different forms of fairness and discrimination.
7. Modeling fairness as a causal hierarchy
Systems thinking allows us to combine three previously independent analytical techniques: machine learning, causal inference, and system dynamics. Doing so allows us to capture the emergent behavior from three different aspects of the fairness problem. We can think of causal Bayesian networks as a higher level representation of supervised machine learning models, which emerges when we “think outside of the black box.” We can think of causal loop diagrams as a higher level representation of causal Bayesian networks, which emerges when we violate the directed acyclic graph constraint and create a feedback loop.
Layer 1. Association
Activity:
Seeing, observing
Layer 2. Intervention
Activity:
Doing, intervening
Layer 3. Counterfactuals
Activity:
Imagining, retrospection, understanding
Layer 1. Association
Field:
Supervised machine learning
Representation: Block diagrams Application: Prediction
Layer 2. Acyclic causation
Field:
Causal inference (Judea Pearl)
Representation: Causal Bayesian networks Application: Control
Layer 3. Cyclic causation
Field:
System dynamics (Jay Forrester)
Representation: Causal loop diagrams Application: Simulation
[…]
Figure 8. Causal hierarchy representation of algorithmic fairness
[…]
Adding a system dynamics layer to Pearl’s causal hierarchy.
Each of the layers in this new causal hierarchy have unique properties and behaviors that make them optimized for different applications. The first layer is associated with supervised machine learning and is optimized for prediction. The second layer is associated with causal inference and is optimized for control. The third layer is associated with system dynamics and is optimized for simulation. These three new causal layers make up an expanded causal hierarchy that is consistent with systems thinking principles. Appendix D shows how this can be part of a larger systems map.
8. Conclusion
This approach allow us to develop a more unified and holistic view of the fairness and machine learning problem, providing essential connections between computer science, politics, and the law. This is the key to building AI systems that are aligned with society’s democratic values and the policymakers’ political agendas. It will enable businesses and regulators to communicate with each other using a common causal language, one that can be used by both data scientists and lawyers. These models may even aid in AI deregulation by identifying mathematical flaws in the existing laws and regulations. This will help businesses to comply with antidiscrimination laws and regulations, reducing business uncertainty and expensive litigation while providing a much needed layer of transparency and trust for consumers of AI products and services.
Systems thinking supplies the missing mathematical models that are needed to build responsible AI systems. To solve fairness, we need to bring together both the social and technical aspects of what is ultimately a complex systems problem. On the social side, it requires modeling the diversity of perspectives that come from both sides of the political spectrum. On the technical side, it requires unifying analytical techniques that were previously thought of as being independent. By unifying all of these concepts together, we hope that this enables the development of AI policies and systems that could lead to practical, real-world solutions for reducing disparities across protected groups in society.
Appendix A Case study: Racial disparities in the United States
We can use the causal models presented in the main part of this paper to demonstrate how to model the complex political debate over racial disparities in the US. According to the US Census Bureau, median household income in 2021 was \$100,573 for Asians, \$74,932 for Whites, \$57,671 for Hispanics, and \$48,297 for Blacks(Survey, 2021). Meanwhile, median household wealth in 2021 was \$320,900 for Asians, \$250,400 for Whites, \$48,700 for Hispanics, and \$27,100 for Blacks(Kochnar and Moslimani, 2023). We would like to understand why Whites and Asians have higher median household income and wealth than Blacks and Hispanics.
[…]
Figure 9. Overt discrimination
[…]
Overt discrimination causes bias in both the decision $D$ and the outcome $Y$. The outcome $Y$ becomes future data $X$ due to a feedback loop. Figure 10 shows what this feedback loop likes within a CBN. We temporarily violate the directed acyclic graph constraint of CBNs and create a path $Y\rightarrow W$, causing the mediator $W$ to become biased.
[…]
Figure 10. Feedback from overt discrimination
[…]
A CBN with a feedback loop from the outcome $Y$ to the mediator $W$.
[…]
Figure 11. Covert discrimination
[…]
So let’s say that US policymakers (particularly Democrats) succeed in instituting affirmative action as shown in Figure 12. While affirmative action might help correct for biases in the decision $D$, notice in Figure 12 that it does not necessarily correct for biases in the outcome $Y$. The protected attribute $A$ does not have a direct effect on the outcome $Y$. It can only directly affect the decision $D$. Correcting for biases in the decision $D$ cannot fully offset biases in the mediator $W$, which means that the outcome $Y$ remains biased. The path $D\rightarrow Y$ only has a weak effect on the outcome $Y$. On the other hand, the path $W\rightarrow Y$ has a strong effect on the outcome $Y$. This is shown in Figure 13.
[…]
Figure 12. Fairness through affirmative action
[…]
Figure 13. Strong and weak effects
[…]
Figure 14. Internal solution and external intervention
[…]
This creates a major problem. Affirmative action may create a false illusion that you can have a system of double standards and not pay the price. While it could enable less qualified Blacks and Hispanics to get access to the some of the same resources that would have went to more qualified Whites and Asians, it doesn’t necessarily translate into better overall outcomes. In college admissions for example, one could argue that a Black or Hispanic college student struggling at an elite university might have been better off thriving at a flagship state university instead. In employment, a Black or Hispanic employee might have been better off being placed in a team where they were at the same level as their White and Asian peers versus being placed in a higher performing team where they ended up dragging the rest of the team down. This is the main argument that has been advanced by Sowell, Mac Donald, and other conservative scholars(Sowell, 2004, 2019; Donald, 2023).
Asians in the US have not historically benefited from affirmative action policies. Asians have instead relied on fairness through unawareness, as shown in Figure 15. In some cases like college admissions, they may even have to outperform their White counterparts to receive the same admission decisions(Grossman et al., 2023). Asians as a group have tended to focus their efforts on the internal solution of competing based on their merits and qualifications. By working against Asians, affirmative action programs may have actually caused Asians to work even harder than other racial groups. By focusing on developing their own human capital instead of demanding strong government intervention, not only have Asians closed the gap in income and wealth with Whites over the long-term, they have actually surpassed them. Over time, this has also reduced perceived biases in the data between Asians and Whites.
[…]
Figure 15. Fairness through unawareness
[…]
On the other hand, some US policymakers (particularly Democrats) have argued for a doubling down on affirmative action and DEI programs. This was especially the case during the past four years under the Biden administration. To address the fact that affirmative action only has a weak effect on outcomes, their goal was to deploy affirmative action and DEI programs across the entire US economy. This would create a racially segregated economic system that provides special treatment for certain minority groups. By instituting a system of double standards and quotas across the US economy, their hope was to equalize outcomes across the racial groups. Doing so would have required a radical restructuring of the US economy, from one that primarily relies on capitalism to one that primarily relies on segregation. It would have reduced the size of the overall economic pie as opportunities are distributed not primarily based on qualifications but instead to reach diversity quotas. As we saw in the recent election, there was not enough popular support in the US for such a radical transformation of the US economy.
These models provide a basis for acknowledging that while historical policies like slavery and Jim Crow generated the initial racial disparities between Whites and Blacks, contemporary policies like affirmative action and DEI may have reinforced those disparities. Affirmative action should at most be used as a temporary solution for marginally closing the racial equity gap. If it is used, then there should be a long-term plan to eventually switch back to fairness through unawareness. If an addictive dependency on affirmative action develops that diverts attention away from internal solutions for closing the merit and qualifications gap, then the affirmative action policy should be ended. Systems thinking shows us that AI researchers should not dismiss fairness through unawareness as a fundamental solution for closing the racial equity gap.
Appendix B Normative effects of fairness and discrimination
Table 4. Normative effects of fairness
Effect Lottery Affirmative action Unaware- ness Suprema- cism
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ – + 0 +
$TE_{a}(d)$ 0 0 0 0
[…]
A list of the normative effects for each form of fairness.
Table 5. Normative effects of discrimination
Effect “No” Reverse Covert Overt
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ + 0 0
$TE_{a}(d)$ + +
[…]
A list of the normative effects for each form of discrimination.
Appendix C Positive effects of fairness and discrimination

### v5

@@ Abstract.
Systems thinking provides us with a way to model the algorithmic fairness problem by allowing us to encode prior knowledge and assumptions about where we believe bias might exist in the data generating process. We can then encode these beliefs as a series of causal graphs, enabling us to link AI/ML systems to politics and the law. This allows us to combine techniques from machine learning, causal inference, and system dynamics in order to capture different emergent aspects of the fairness problem. We can use systems thinking to help policymakers on both sides of the political aisle to understand the complex trade-offs that exist from different types of fairness policies, providing a sociotechnical foundation for designing AI policy that is aligned to their political agendas and with society’s shared democratic values.
@@ 1. Introduction
For several years, AI researchers have struggled with how to build AI/ML systems that are fair and aligned to society’s shared democratic values. Purely technical approaches that rely on correlation alone have fallen far short of what is needed to solve this difficult problem. Systems thinking offers a novel approach to the algorithmic fairness problem by allowing us to use causality to combine complex ideas from the social sciences and computer science into a comprehensive and integrated theory. By leveraging prior knowledge and assumptions about where we believe bias might exist in the data generating process, we can help policymakers on both sides of the aisle to design AI policy that is aligned to their political agendas.
@@
We will then turn our attention to the technical aspects of the fairness problem, which forms the bulk of this paper. We will use systems thinking to show how to combine analytical techniques from machine learning, causal inference, and system dynamics to model different facets of the fairness problem. Machine learning is necessary for building AI systems that can make fair predictions. Causal inference allows us to visualize different forms of bias, showing us what interventions are needed to make a machine learning model fair. System dynamics allows us to model bias from the data generating process, providing counterintuitive explanations as to why disparities continue to persist across protected groups. Finally, we will demonstrate how these three techniques can form the basis of a new causal hierarchy for modeling algorithmic fairness.
@@ 2. Related work
Pearl has argued for a few decades that the discrimination problem is fundamentally a causality problem(Pearl, 2001). He has argued that fairness can be modeled using causal Bayesian networks (CBNs) as a series of direct, indirect, and spurious effects. This work was later expanded upon by Plecko and Bareinboim, who proposed the idea of a standard fairness model(Plecko and Bareinboim, 2022). They argued that there exists a causal template through which entire classes of fairness models can be derived using a small number of modeling assumptions. Kusner et al. argued for the use of counterfactuals in modeling fairness(Kusner et al., 2017). Chiappa and Isaac showed that certain paths within a CBN could be labeled as fair or unfair(Chiappa and Isaac, 2019). Further attempts at building causal fairness models were made by Nilforoshan et al.(Nilforoshan et al., 2022) and Kilbertus et al.(Kilbertus et al., 2017). Barocas et al. devoted a chapter in their textbook on fairness and machine learning to discuss how causality might be used to model fairness(Barocas et al., 2023).
However, all of their approaches were limited due to their inability to construct a valid causal model of the problem. This is because their approaches did not fully take into account the social and political aspects of the fairness problem. Society has different ideological beliefs about what is fair and unfair based on different sets of assumptions about where bias is within the data generating process. Our paper addresses this issue by encoding society’s different mental models of bias into causal models using first principles. By taking into account a diversity of perspectives, we can understand the normative aspects of the fairness problem. This is necessary for linking machine learning to politics and ultimately to the law.
Another key difference is that their causal approaches relied solely on causal inference, which has limitations in its ability to model the dynamics of the data generating process. Our approach leverages systems thinking and system dynamics to gain a fuller understanding of the data generating process through the modeling of feedback loops. Martin has also explored modeling AI bias using community-based system dynamics(Jr, 2023, 2021). Furthermore, we link supervised machine learning to causal inference by recognizing that a CBN can be used as a higher level abstraction of a black box ML model. By bringing supervised machine learning, causal inference, and systems thinking together, we can model different aspects of the algorithmic fairness problem.
@@ 3. Towards a unified theory of fairness
To determine whether an outcome is biased and unfair, all we need to know is whether we would attribute a bad outcome as being internal or external. This makes algorithmic fairness a type of credit assignment problem. But since we usually focus on bad outcomes, this oftentimes becomes a blame assignment problem. That being said, identifying attribution is much harder in practice, as we will show later when we model fairness using system dynamics. Both internal and external factors may influence and reinforce each other, leading to vicious feedback loops. This becomes a particularly difficult problem when trying to explain disparities across protected groups like race, which we describe in greater detail in Section 6.2.
This causality dilemma between attributing disparities to internal or external factors ultimately leads to the controversial and heated political debate over how best to reduce disparities across protected groups(Sowell, 2019; Donald, 2023; Mills, 1997; Kendi, 2017; Loury, 2002). Those on the political left are more likely to attribute disparities as being due to external factors, which provides the moral justification for strong government interventions like affirmative action or DEI. The left also tends to focus more on legal forms of discrimination like disparate impact or indirect discrimination. Those on the political right are more likely to attribute disparities as being due to internal factors, which provides the moral justification for limited government intervention like fairness through unawareness. The right also tends to focus more on legal forms of discrimination like disparate treatment or direct discrimination. This ideological debate is fundamental to understanding algorithmic fairness, but has been largely missing from the academic literature.
@@ 5. Modeling fairness as a causal inference problem
Note that this is not an exhaustive list of all forms of fairness and discrimination. For example, we could have a form of covert discrimination where the edges of Figure 2d are all colored in blue instead of red.
@@ 5.1. Mapping fairness to politics
Left-wing notions of fairness rely on affirmative action, which is associated with the criterion $D\perp\!\!\!\perp A$. That is, the decision $D$ is independent of the protected attribute $A$, which is achieved by using the protected attribute $A$ to correct for biases in the data $X$ that is used in the decision $D$. We imagine what the applicant’s decision would have been had they been a member of a counterfactual class. We then perform an intervention to obtain the same decision. For example, a college admissions decision would affirmatively use race to correct for deficiencies in the applicant’s merit due to systemic racism. This leads to group fairness, where different groups receive more similar aggregate decision rates(Castelnovo et al., 2022; Verma and Rubin, 2018; Mehrabi et al., 2021; Binns, 2019). This form of fairness relies on Pearl’s first fundamental law of causal inference: counterfactuals and interventions.
@@ 5.2. Mapping discrimination to antidiscrimination law
When a protected attribute $A$ has a direct effect on a decision $D$ ($A\rightarrow D$), we might consider this as a form of disparate treatment or direct discrimination. This could map to overt discrimination (Figure 2b) and reverse discrimination (Figure 2f). To avoid disparate treatment or direct discrimination, the protected attribute $A$ should have no direct effect on the decision $D$.
@@
When a protected attribute $A$ has an indirect effect on a decision $D$, we might consider this as a form of disparate impact or indirect discrimination. This could map to covert discrimination (Figure 2d) and “no” discrimination (Figure 2h). Even though the protected attribute $A$ has no direct effect on a decision $D$ (making the policy neutral), the inclusion of biased data or exclusion of unbiased data could cause the protected attribute $A$ to have an indirect effect on the decision $D$ that adversely affects a protected group.
For disparate impact in the US, there is a three-step test(of Justice Civil Rights Division, 2024). The first step is to establish that a neutral policy has an adverse disparate impact on a protected group. This would involve testing whether there is a disparity in decisions $D$ across the protected class $A$. The second step is to establish the business necessity of that policy. This would involve arguing that the data $X$ being used has some causal relationship to the mediator $W$. The final step is to identify a less discriminatory alternative model, which would involve performing a search for a counterfactual model.
CBNs provide an analytical framework to reason about whether a model is causing illegal discrimination. By being able to explicitly model the different forms of discrimination, CBNs provide a common language that businesses and regulators can use to ensure that machine learning models are compliant with laws and regulations.
5.3. Normative effects of fairness and discrimination
Table 4. Normative effects of fairness
Effect Lottery Affirmative action Unaware- ness Suprema- cism
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ – + 0 +
$TE_{a}(d)$ 0 0 0 0
[…]
A list of the normative effects for each form of fairness.
Table 5. Normative effects of discrimination
Effect “No” Reverse Covert Overt
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ + 0 0
$TE_{a}(d)$ + +
[…]
A list of the normative effects for each form of discrimination.
@@ 6. Modeling fairness as a system dynamics problem
Finally, we add in the “shifting the burden / addiction” archetype. If one group is allocated less resources than another group, then it could breed resentment by the disadvantaged group. That group could then demand that the protected attribute $A$ be used to externally intervene on the allocation decision $D$. This external intervention is equivalent to fairness through affirmative action (Figure 2e). While it might reduce allocation disparities in the short term, it may actually create a dependence on the external intervention over the long term. This is because if allocation decisions $D$ are not based purely on data $X$, which is a measure of mediator $W$ (e.g. merit, qualifications), then the assignment of resources to the disadvantaged group may not lead to their success $Y$. For example, approving a loan to a minority applicant even though the applicant is not creditworthy could lead to a worse outcome (default) than if the loan was denied in the first place. That decreased success in outcome $Y$ for the less successful group could trigger a demand for an even stronger external intervention $A$ on the decision $D$, reinforcing this problem. In the meanwhile, attention may be taken away from the internal solution of optimizing for the mediator $W$ and the data $X$ so that the disadvantaged group would perform better under fairness through unawareness. This may especially be the case if the disadvantaged group believes that the mediator $W$ is biased and therefore does not focus on optimizing that metric.
6.1. Positive effects of fairness and discrimination